OpenStudy (anonymous):

plz i need help :( Evaluate the expression if a = 1 and b = 2. fraction numerator 5 left parenthesis a star times plus 3 b right parenthesis over denominator a b end fraction A. 20 B. 5.5 C. 15 D. 17.5

OpenStudy (freckles):

\[\frac{5( \star \cdot 3b)}{ab}\] any idea what the value we should plug in for the star ?

OpenStudy (anonymous):

1

OpenStudy (freckles):

so you are given \[a=1, b=2, \star=1\]

OpenStudy (freckles):

replace a with 1 replace b with 2 replace star with 1 then \[\frac{5( \star \cdot 3b)}{ab}=\frac{5(1 \cdot 3(2))}{1(2)}\]

OpenStudy (anonymous):

no the star is a

OpenStudy (freckles):

oh your problem said something about a star in it

OpenStudy (anonymous):

and a is 1

OpenStudy (anonymous):

sorry about that

OpenStudy (freckles):

I'm sorry kind of confused about times plus and what that means exactly

OpenStudy (freckles):

can you draw the problem or is it really worded like that?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):